# NovelSeek: When Agent Becomes the Scientist – Building Closed-Loop System from Hypothesis to Verification

### v2

@@ 3.3 Insightful Analyses
Improving Baseline in Multi-Dimension. NovelSeek not only can improve the performance on different tasks, but it also enhance the quality of ideas in other dimensions. For example, as shown in Tab. 7, on few-shot yield prediction task. We find that the results of baseline methods are unstable, for the results of multiple repeated experiments tend to exhibit large variance (e.g., 24.2$\pm$4.2 when train-set=60). In contrast, the methods proposed by NovelSeek can improve both the performance and the stability of the results. For example, when the train-set=60, ADAGT proposed by NovelSeek achieves 34.8 average $R^{2}$ across 5 repeated experiments compared to 24.2 achieved by the baseline method. Besides, the variance of the results achieved by ADAGT (i.e., ±$\pm 1.1$) is much lower than baseline methods (i.e., ±$\pm 4.2$). This phenomenon further shows the quality of ideas and code implementation of NovelSeek.
Comparison with AI-Researcher. We evaluate the performance and cost of NovelSeek and AI-Researcher  (Lab, 2025) on AutoRYP and Auto2DCls research tasks. To ensure a fair comparison, we supplied AI-Researcher with the same code templates that NovelSeek uses. Both systems employ GPT-4o-2024-08-06 for idea generation and Claude-3-7-Sonnet-20250219 for code generation. As shown in Tab. 9, NovelSeek outperforms both the baseline methods and AI-Researcher across both tasks, whereas AI-Researcher is unable to improve the provided baselines. One important reason for NovelSeek’s outstanding performance is its ability to generate novel ideas through extensive search and reflection. Moreover, NovelSeek has a complete experimental planning and adaptive evolution mechanism, thereby enabling it to achieve better performance. In contrast, the idea generated by AI-Researcher is more dependent on user-provided reference papers, limiting its novelty. Moreover, AI-Researcher often ignore the prior information of the existing codebases, which further hinders its performance. In terms of cost, NovelSeek is significantly more economical than AI-Researcher. For an instance, the economic cost required for NovelSeek is approximately one-sixth that of AI-Researcher on the AutoRYP task. This lower cost allows NovelSeek to conduct broader scientific experiments, thereby accelerating the exploration and validation of innovative research ideas.
@@ 4.1 Qualitative Results
In this section, we present the results from various tasks implemented using NovelSeek.
@@ 4.1.1 Visual Examples of Various Tasks
We present showcases for three distinct tasks: AutoRYP, AutoMD, and AutoPower, to highlight the innovative methodologies discovered and their applications. These showcases are illustrated in Figs.  5,  6, and  9, respectively. Each task demonstrates a unique approach to solving complex problems, showcasing the potential impact of NovelSeek across different scientific domains.
In the AutoRYP task illustrated in Fig. 5, NovelSeek autonomously discovered an innovative approach called "Adaptive Dual-Attention Graph-Transformer with Dynamic Freezing" for predicting chemical reaction yields. This method effectively integrates SMILES and graph-derived descriptors using a hybrid graph-transformer network, incorporating hierarchical attention mechanisms to enhance accuracy while minimizing overfitting. The approach features a Dual-Attention Fusion Mechanism (DAFM) that systematically combines token and graph embeddings with reaction conditions, ensuring effective information flow across different representations. Furthermore, a dynamic layer freezing mechanism, based on gradient magnitudes, optimizes which layers are trained, thereby enhancing generalization in low-data scenarios. Implemented with self-attention and cross-modality attention modules, this system not only combines 1D and 2D molecular representations effectively but also improves prediction accuracy and model adaptability. Consequently, this showcase underscores the method’s potential for advancing research in complex chemical tasks using deep learning.
@@ 4.2 Human Evaluation
Table 10 compares the novelty of ideas generated by our NovelSeek and AI-Scientist-V2 (Yamada et al., 2025), across various research tasks. Each task involves generating 20 ideas, which are evaluated by five qualified reviewers. The assessments focus on four criteria: soundness, contribution, overall rating, and confidence. For each research task, the average scores of the 20 ideas are reported.
In the Reaction Yield Prediction task, NovelSeek outperforms AI-Scientist-V2 in all aspects, especially in overall rating and soundness. Similarly, for 2D Semantic Segmentation, NovelSeek shows better idea generation ability, particularly in soundness and overall rating. In 2D Image Classification and Point Cloud Autonomous Driving, NovelSeek scores higher across all criteria, indicating a consistent advantage over AI-Scientist-V2 in generating more effective and novel ideas.
Table 10: From the perspectives of soundness, contribution, and overall, we compare the novelty of ideas generated by NovelSeek and AI-Scientist-V2 (Yamada et al., 2025). For each research task, we generate 20 ideas. Each idea is scored by 5 qualified reviewers, and the final score for each task is reported as the average score of all 20 ideas. The detailed scores for each idea can be found in the Appendix B.1.
@@
Reaction Yield Prediction NovelSeek 3.09 2.66 4.35 4.00
@@
2D Semantic Segmentation NovelSeek 2.41 2.35 4.05 3.48
@@
2D Image Classification NovelSeek 3.15 3.10 5.85 3.32
@@
Point Cloud Autonomous Driving NovelSeek 2.75 2.95 5.10 4.10
@@ 6 Conclusion and Future Works
Summary. We have introduced a closed-loop multi-agent framework for the first time, which supports 12 types of scientific research tasks. It has been validated to generate novel ideas and transform these ideas into code that can effectively improve performance. NovelSeek refines the initially generated ideas through human-interactive feedback enriched with a self-evolutionary path of ideas. It facilitates the transformation from coarse-grained proposals to fine-grained methodologies via an idea-to-methodology construction process. Furthermore, by leveraging multi-round experimental planning and execution, it implements the corresponding theoretical methods, thereby completing the closed-loop process in scientific research—from hypothesis generation to verification.
Future Outlook. NovelSeek faces several key technical challenges that need to be addressed in the future:
@@
  • Scientific Discovery-related Benchmark Construction: This involves evaluating the value that an idea can bring, rather than simply evaluating its novelty. It also includes evaluating whether the methods proposed by AI align with their corresponding code implementations and determining whether NovelSeek demonstrates a certain level of generalization ability in broader scientific scenarios.
@@ 1. Core Contributors
2. Multi-agent System Algorithm Design
Self-Evolving Idea: Shiyang Feng
Idea-to-Methodology: Xiangchao Yan
Coder: Shiyang Feng, Jiakang Yuan, Xiangchao Yan
@@ B.1 Scoring Criteria for Idea Review
In Table 10 of the main text, we conducted a human evaluation to assess the novelty of ideas generated by AI-Scientist-V2 (Yamada et al., 2025) and our NovelSeek. The evaluation was carried out across four dimensions: soundness, contribution, overall rating, and confidence.
Specifically, considering the evaluation cost, we opted to evaluate four types of research tasks: reaction yield prediction, 2D semantic segmentation, 2D image classification, and point cloud autonomous driving. Each invited researcher was required to have peer-review qualifications for top-tier journals or conferences in the relevant field. For each research task, we generated 20 ideas using both AI-Scientist-V2 and NovelSeek, and five experienced researchers were invited to score each idea. In this part, we provide a detailed description of the scoring criteria for each reviewer, as outlined below:
@@ B.2 Information on the Expert Review Process
  • Each evaluator is assigned 20 ideas generated by Novelseek and 20 ideas generated by AI-Scientist-V2 (Yamada et al., 2025). For each idea, evaluators must carefully review the generated content and provide final scores across four dimensions: Soundness, Contribution, Overall, and Confidence.
@@
Figure 9: Visual Examples of AutoPower Task
Figure 10: Visual Examples of AutoSenCls Task
Figure 11: Visual Examples of Auto3DCls Task
Figure 12: Visual Examples of Auto2DCls Task
Appendix C NovelSeek Software Development
Figure 13: NovelSeek software platform includes features such as the user entry interface, task selection interface, idea-tree visualization and human-computer interaction interface, code generation, and auto-debug interface. In the near future, we plan to support additional functionalities, including custom dataset uploads and academic idea thinking modes.
Fig. 13 shows the front-end interface of the current NovelSeek software platform. Overall, NovelSeek software platform employs a frontend-backend separation design pattern, building a highly scalable distributed service platform.
@@ Appendix C NovelSeek Software Development
Appendix D Visualization Results
We further conducted detailed visualization and analysis around the multi-round experimental planning and execution (Fig. 8) in NovelSeek, as well as the automated scientific research tasks (Figs. 91011 and 12) supported by NovelSeek.
Fig. 8 illustrates the process of Experimental Planning and Adaptive Evolution on AutoTPPR. Each block in the figure represents a step in the multi-round experiment planning process, where the complete NovelSeek-generated method is decomposed into multiple logical steps. This allows for task decomposition during the experimental validation phase, thereby facilitating more significant benchmark results.

### v3

@@ 3.3 Insightful Analyses
Improving Baseline in Multi-Dimension. InternAgent not only can improve the performance on different tasks, but it also enhance the quality of ideas in other dimensions. For example, as shown in Tab. 7, on few-shot yield prediction task. We find that the results of baseline methods are unstable, for the results of multiple repeated experiments tend to exhibit large variance (e.g., 24.2$\pm$4.2 when train-set=60). In contrast, the methods proposed by InternAgent can improve both the performance and the stability of the results. For example, when the train-set=60, ADAGT proposed by InternAgent achieves 34.8 average $R^{2}$ across 5 repeated experiments compared to 24.2 achieved by the baseline method. Besides, the variance of the results achieved by ADAGT (i.e., ±$\pm 1.1$) is much lower than baseline methods (i.e., ±$\pm 4.2$). This phenomenon further shows the quality of ideas and code implementation of InternAgent.
Comparison with AI-Researcher. We evaluate the performance and cost of InternAgent and AI-Researcher  (Lab, 2025) on AutoRYP and Auto2DCls research tasks. To ensure a fair comparison, we supplied AI-Researcher with the same code templates that InternAgent uses. Both systems employ GPT-4o-2024-08-06 for idea generation and Claude-3-7-Sonnet-20250219 for code generation. As shown in Tab. 9, InternAgent outperforms both the baseline methods and AI-Researcher across both tasks, whereas AI-Researcher is unable to improve the provided baselines. One important reason for InternAgent’s outstanding performance is its ability to generate novel ideas through extensive search and reflection. Moreover, InternAgent has a complete experimental planning and adaptive evolution mechanism, thereby enabling it to achieve better performance. In contrast, the idea generated by AI-Researcher is more dependent on user-provided reference papers, limiting its novelty. Moreover, AI-Researcher often ignore the prior information of the existing codebases, which further hinders its performance. In terms of cost, InternAgent is significantly more economical than AI-Researcher. For an instance, the economic cost required for InternAgent is approximately one-sixth that of AI-Researcher on the AutoRYP task. This lower cost allows InternAgent to conduct broader scientific experiments, thereby accelerating the exploration and validation of innovative research ideas.
@@ 4.1 Qualitative Results
In this section, we present the results from various tasks implemented using InternAgent.
@@ 4.1.1 Visual Examples of Various Tasks
We present showcases for three distinct tasks: AutoRYP, AutoMD, and AutoPower, to highlight the innovative methodologies discovered and their applications. These showcases are illustrated in Figs.  5,  6, and  9, respectively. Each task demonstrates a unique approach to solving complex problems, showcasing the potential impact of InternAgent across different scientific domains.
In the AutoRYP task illustrated in Fig. 5, InternAgent autonomously discovered an innovative approach called "Adaptive Dual-Attention Graph-Transformer with Dynamic Freezing" for predicting chemical reaction yields. This method effectively integrates SMILES and graph-derived descriptors using a hybrid graph-transformer network, incorporating hierarchical attention mechanisms to enhance accuracy while minimizing overfitting. The approach features a Dual-Attention Fusion Mechanism (DAFM) that systematically combines token and graph embeddings with reaction conditions, ensuring effective information flow across different representations. Furthermore, a dynamic layer freezing mechanism, based on gradient magnitudes, optimizes which layers are trained, thereby enhancing generalization in low-data scenarios. Implemented with self-attention and cross-modality attention modules, this system not only combines 1D and 2D molecular representations effectively but also improves prediction accuracy and model adaptability. Consequently, this showcase underscores the method’s potential for advancing research in complex chemical tasks using deep learning.
@@ 4.2 Human Evaluation
Table 10 compares the novelty of ideas generated by our InternAgent and AI-Scientist-V2 (Yamada et al., 2025), across various research tasks. Each task involves generating 20 ideas, which are evaluated by five qualified reviewers. The assessments focus on four criteria: soundness, contribution, overall rating, and confidence. For each research task, the average scores of the 20 ideas are reported.
In the Reaction Yield Prediction task, InternAgent outperforms AI-Scientist-V2 in all aspects, especially in overall rating and soundness. Similarly, for 2D Semantic Segmentation, InternAgent shows better idea generation ability, particularly in soundness and overall rating. In 2D Image Classification and Point Cloud Autonomous Driving, InternAgent scores higher across all criteria, indicating a consistent advantage over AI-Scientist-V2 in generating more effective and novel ideas.
Table 10: From the perspectives of soundness, contribution, and overall, we compare the novelty of ideas generated by InternAgent and AI-Scientist-V2 (Yamada et al., 2025). For each research task, we generate 20 ideas. Each idea is scored by 5 qualified reviewers, and the final score for each task is reported as the average score of all 20 ideas. The detailed scores for each idea can be found in the Appendix B.1.
@@
Reaction Yield Prediction InternAgent 3.09 2.66 4.35 4.00
@@
2D Semantic Segmentation InternAgent 2.41 2.35 4.05 3.48
@@
2D Image Classification InternAgent 3.15 3.10 5.85 3.32
@@
Point Cloud Autonomous Driving InternAgent 2.75 2.95 5.10 4.10
@@ 6 Conclusion and Future Works
Summary. We have introduced a closed-loop multi-agent framework for the first time, which supports 12 types of scientific research tasks. It has been validated to generate novel ideas and transform these ideas into code that can effectively improve performance. InternAgent refines the initially generated ideas through human-interactive feedback enriched with a self-evolutionary path of ideas. It facilitates the transformation from coarse-grained proposals to fine-grained methodologies via an idea-to-methodology construction process. Furthermore, by leveraging multi-round experimental planning and execution, it implements the corresponding theoretical methods, thereby completing the closed-loop process in scientific research—from hypothesis generation to verification.
Future Outlook. InternAgent faces several key technical challenges that need to be addressed in the future:
@@
  • Scientific Discovery-related Benchmark Construction: This involves evaluating the value that an idea can bring, rather than simply evaluating its novelty. It also includes evaluating whether the methods proposed by AI align with their corresponding code implementations and determining whether InternAgent demonstrates a certain level of generalization ability in broader scientific scenarios.
@@ 1. Core Contributors
2. General Multi-agent System
Deep Research Agent: Yusong Hu, Runmin Ma, Bo Zhang
Self-Evolving Idea Agent: Shiyang Feng
Idea-to-Methodology Agent: Xiangchao Yan
Coding Agent: Shiyang Feng, Jiakang Yuan, Xiangchao Yan
@@ B.1 Scoring Criteria for Idea Review
In Table 10 of the main text, we conducted a human evaluation to assess the novelty of ideas generated by AI-Scientist-V2 (Yamada et al., 2025) and our InternAgent. The evaluation was carried out across four dimensions: soundness, contribution, overall rating, and confidence.
Specifically, considering the evaluation cost, we opted to evaluate four types of research tasks: reaction yield prediction, 2D semantic segmentation, 2D image classification, and point cloud autonomous driving. Each invited researcher was required to have peer-review qualifications for top-tier journals or conferences in the relevant field. For each research task, we generated 20 ideas using both AI-Scientist-V2 and InternAgent, and five experienced researchers were invited to score each idea. In this part, we provide a detailed description of the scoring criteria for each reviewer, as outlined below:
@@ B.2 Information on the Expert Review Process
  • Each evaluator is assigned 20 ideas generated by InternAgent and 20 ideas generated by AI-Scientist-V2 (Yamada et al., 2025). For each idea, evaluators must carefully review the generated content and provide final scores across four dimensions: Soundness, Contribution, Overall, and Confidence.
@@
Figure 9: Result Analysis (Generated Idea, Method, and Code) of AutoPower Task
Figure 10: Result Analysis (Generated Idea, Method, and Code) of AutoSenCls Task
Figure 11: Result Analysis (Generated Idea, Method, and Code) of Auto3DCls Task
Figure 12: Result Analysis (Generated Idea, Method, and Code) of Auto2DCls Task
Appendix C InternAgent Software Development and Application Interface
Figure 13: InternAgent software platform includes features such as the user entry interface, task selection interface, idea-tree visualization and human-computer interaction interface, code generation, and auto-debug interface. In the near future, we plan to support additional functionalities, including custom dataset uploads and academic idea thinking modes.
Software Development. Fig. 13 shows the front-end interface of the current InternAgent software platform. Overall, InternAgent software platform employs a frontend-backend separation design pattern, building a highly scalable distributed service platform.
@@ Appendix C InternAgent Software Development and Application Interface
InternAgent Trial Application: Accelerating Scientific Research. InternAgent is the first end-to-end Autonomous Scientific Research (ASR) multi-agent framework. The release of this innovative achievement enables Agent to autonomously analyze problems, review literature, conduct deep research, reflect, and carry out experiments, much like human scientists. It can automatically complete the entire process from hypothesis generation to experimental validation, starting from an initial research idea. This marks the official arrival of a new era in AI-driven scientific discovery.
Due to InternAgent’s reliance on massive API key calls and GPU compute clusters for executing experiments, there are associated costs. As a result, we have recently launched a whitelist application page. On this page, applicants are required to provide information about the scenarios in which they will use InternAgent, their research tasks, and the institution they belong to, so that we can select the initial batch of seed users in InternAgent. The link to the application experience page is as follows: https://discovery.intern-ai.org.cn/.
Appendix D Result Analysis
We further conducted detailed visualization and analysis around the multi-round experimental planning and execution (Fig. 8) in InternAgent, as well as the automated scientific research tasks (Figs. 91011 and 12) supported by InternAgent.
Fig. 8 illustrates the process of Experimental Planning and Adaptive Evolution on AutoTPPR. Each block in the figure represents a step in the multi-round experiment planning process, where the complete InternAgent-generated method is decomposed into multiple logical steps. This allows for task decomposition during the experimental validation phase, thereby facilitating more significant benchmark results.